OpenStudy (elleblythe):

Evaluate: sin(8pi/15)cos(pi/5) - cos(8pi/15)sin(pi/5) Show solutions please. This isn't graded or anything, just doing some practice exercises so I can understand the lessons better.

OpenStudy (anonymous):

its in the form of sinAcosB-cosAsinB=sin(A-B) here A=8pi/15 B=pi/5 (A-B)=(8pi/15)-(pi/5) =pi/3=60degrees so sin(A-B)=sin(pi/3) final answer will be \[\sqrt{3}\div2\]